Laker Incorporated’s fiscal year-end is December 31, 2021. The following is an adjusted trial balance as of December 31.  

Calculate the ending balance of Retained Earnings :
  

Accounts Debit Credit
Cash $ 11,700
Supplies 37,500
Prepaid Rent 28,500
Accounts Payable $ 2,700
Notes Payable 27,000
Common Stock 38,500
Retained Earnings 8,700
Dividends 3,700
Service Revenue 54,300
Salaries Expense 19,700
Advertising Expense 12,700
Rent Expense 9,700
Utilities Expense 7,700
Totals $ 131,200 $ 131,200
Answer #1

Retained Earnings formula

RE=Beginning period retained earnings +Net income/loss -Dividends

Given

Beginning retained earnings =8,700

Revenue = 54300

Dividends=3,700

We should calculate Net income

Net income=Revenue -Total expenses

Total Expenses =Salary Expense+ Advertising expenses+ Rent Expenses + utilities expenses

=19700+12700+9700+7700 =49800

Net income =54300-49800

              = 4500

Now let us calculate RE

RE=Beginning period retained earnings +Net income/loss -Dividends

=8700+4500-3700

=9,500